### v2

# Gaga: Group Any Gaussians via 3D-aware Memory Bank

Abstract
We introduce Gaga, a framework that reconstructs and segments open-world 3D scenes by leveraging inconsistent 2D masks predicted by zero-shot class-agnostic segmentation models. Contrasted to prior 3D scene segmentation approaches that rely on video object tracking or contrastive learning methods, Gaga utilizes spatial information and effectively associates object masks across diverse camera poses through a novel 3D-aware memory bank. By eliminating the assumption of continuous view changes in training images, Gaga demonstrates robustness to variations in camera poses, particularly beneficial for sparsely sampled images, ensuring precise mask label consistency. Furthermore, Gaga accommodates 2D segmentation masks from diverse sources and demonstrates robust performance with different open-world zero-shot class-agnostic segmentation models, significantly enhancing its versatility. Extensive qualitative and quantitative evaluations demonstrate that Gaga performs favorably against state-of-the-art methods, emphasizing its potential for real-world applications such as 3D scene understanding and manipulation.
Figure 1: Gaga groups any Gaussians in an open-world 3D scene and renders multi-view consistent segmentation (pixels of the same region across views are represented with the same color). By employing a 3D-aware memory bank, we eliminate the label inconsistency that exists in 2D segmentation predicted by foundational models and assign each mask across different views a universal group ID. This enables the process of lifting 2D segmentation to a consistent 3D segmentation. Gaga produces accurate 3D object segmentation, achieving high-quality results for downstream applications such as scene manipulation (e.g. changing the cushion’s color of the footstool to maroon).
1 Introduction
Figure 2: Comparison of rendered segmentation. Contrastive learning-based methods, such as OmniSeg [33], do not provide unique mask labels to each segmentation group, leading to inconsistencies across multiple views (e.g., the coffee table). Gaussian Grouping [32] addresses multi-view segmentation by utilizing a video tracker, but it often misidentifies objects when similar items are present (e.g., the leather sofa) and struggles with significant camera perspective changes. In contrast, Gaga ensures multi-view consistent segmentation masks, overcoming these limitations.
Effective open-world 3D segmentation is essential for scene understanding and manipulation. Despite notable advancements in 2D open-world segmentation techniques, exemplified by Segment Anything (SAM) [16] and EntitySeg [23], extending these methodologies to the realm of 3D encounters the challenge of ensuring consistent mask label assignment across multi-view images. Specifically, masks of the same object across different views may have different mask label IDs, as the 2D class-agnostic segmentation model independently processes the multi-view images. Naively lifting these inconsistent 2D masks to 3D introduces ambiguity and leads to inferior results in 3D scene segmentation.
Existing works [33, 15, 32, 11, 9] address this issue either by contrastive learning or video object tracking. One line of works build upon 3D scenes represented as 3D Gaussians [13]. In this approach, a feature vector is learned for each 3D Gaussian through a contrastive learning paradigm. The method encourages similar features for 3D Gaussians whose projections fall within the same segmentation mask, while pushing apart those belonging to different segmentation regions. Segmentation maps for each view are then obtained by clustering these feature vectors. However, these methods do not assign a specific mask label to each 3D segmentation group, resulting in inconsistent multi-view segmentation, as shown in Fig. 2. Alternatively, some methods [32, 11] tackle multi-view inconsistency in 3D scene segmentation by treating multi-view images as a video sequence and adopting an off-the-shelf video object tracking method [7]. Nevertheless, this design assumes minimal view changes between multi-view images—a condition that may not always hold in real-world 3D scenes, particularly when input views are sparse. Consequently, these methods often struggle with similar objects or occluded objects that intermittently disappear and reappear in the sequence, as shown in Fig. 2.
In this paper, we identify a fundamental limitation in existing open-world 3D scene segmentation methods that leads to 3D inconsistency: the inadequate exploitation of 3D information inherently provided by the scene. Our main intuition is that masks of the same object across different views shall correspond to the same group of 3D Gaussians. Hence, we can assign identical universal mask IDs to masks from different views when there is a large overlap between their corresponding 3D Gaussian groups.
Based on this intuition, we propose Gaga, which groups any 3D Gaussians and renders consistent 3D class-agnostic segmentation across different views. Given a collection of posed RGB images, we first employ Gaussian Splatting [13] to reconstruct the 3D scene and extract 2D masks using an open-world segmentation method [16, 23]. Subsequently, we iteratively build a 3D-aware memory bank that collects and stores the Gaussians grouped by category. For each input view, we project each 2D mask into 3D space using camera parameters and search the memory bank for the category with the largest overlapping with the unprojected mask. Based on the degree of overlapping, we either assign the mask to an existing category or create a new one. Finally, following the mask association process described above, we can get a set of multi-view consistent 2D segmentation masks. Finally, we learn a feature vector (i.e., identity encoding [32]) for each 3D Gaussian that encodes its category information. Specifically, we splat the feature vectors onto 2D image and decode a segmentation map through a linear layer. The predicted masks are then compared with the segmentation masks obtained from our 3D-aware memory bank for supervision. This approach ensures that our identity encoding remains multi-view consistent, thanks to the consistent segmentation masks across different viewpoints. Our contributions include:
  • We propose a framework that reconstructs and segments 3D scenes using inconsistent 2D masks generated by open-world segmentation models.
  • To resolve the inconsistency of 2D masks across views, we design a 3D-aware memory bank that collects Gaussians of the same semantic group. This memory bank is then employed to align 2D masks across diverse views.
  • We show that the proposed method can effectively leverage any 2D class-agnostic segmentation masks, making it readily applicable for synthesizing novel view images and segmentation masks.
  • Comprehensive experiments on diverse datasets and challenging scenarios, including sparse input views, demonstrate the effectiveness of the proposed method both qualitatively and quantitatively.
2 Related Work
Segment and Tracking Anything in 2D. Segment Anything (SAM) [16] and EntitySeg [23] demonstrate the effectiveness of large-scale training in image segmentation, thus establishing a pivotal foundation for open-world segmentation methods. Subsequent studies [31, 8, 7] further extend the applicability of SAM to video data by leveraging video object segmentation algorithms to propagate SAM masks. Conversely, acquiring data for training their 3D counterparts poses a challenge, given that existing large-scale 3D datasets with annotated segmentation [27, 10] primarily focus on indoor scenarios.
NeRF-based 3D Segmentation. Neural Radiance Fields (NeRFs) [21] model scenes as continuous volumetric functions, learned through neural networks that map 3D coordinates to scene radiance. This approach facilitates the capture of intricate geometric details and the generation of photorealistic renderings, offering novel view synthesis capabilities. Semantic-NeRF [35] initiates the incorporation of semantic information into NeRFs and enables the generation of semantic masks for novel views. Note that semantic segmentation masks do not face the challenge of ambiguous mask ID across views. Numerous methods expand the scope by introducing instance modeling and matching instance masks relying on existing 3D bounding boxes [20, 12], resorting to the cost-based linear assignment during training [26, 29] or directly training instance-specific MLPs [17]. However, most of these methods are developed based on ground truth segmentation and tailored for scene modeling within specific domains. They often entail high computational costs and lack substantial evidence of their performance in open-world scenarios.
Leveraging SAM’s open-world segmentation capability, SA3D [3] endeavors to recover a 3D consistent mask by tracing 2D masks across adjacent views with user guidance. Similarly, Chen et al. [6] distill SAM encoder features into 3D and query the decoder. In contrast, Gaga achieves multi-view consistency without user intervention, offering segmentation for all objects rather than an instance. Garfield [15] densely samples SAM masks and trains a scale-conditioned affinity field supervised on the scale of each mask deprojected to 3D. However, this work focuses on 3D clustering rather than segmentation.
Gaussian-based 3D Segmentation. As an alternative to NeRF and its variants [21, 4, 22, 28], Gaussian Splatting [13, 5, 30, 34] has recently emerged as a powerful approach to reconstruct 3D scenes via real-time radiance field rendering. By representing the scene as 3D Gaussians from posed images, it achieves photorealistic novel view synthesis with high reconstruction quality and efficiency. Additionally, manipulating 3D Gaussians for scene editing is more straightforward compared to NeRF’s representation.
SAGA [2] renders a 2D SAM feature map and uses a SAM guidance loss to learn 3D segmentation from the ambiguous 2D masks. Similar to [3], this method requires user input and only provides segmentation for one object at a time. Feature 3DGS [36] distills LSeg [18] and SAM features to 3D Gaussians and decodes rendered features to obtain segmentation. However, it fails to provide consistent segmentation across views. Gaussian Grouping [32] and CoSSegGaussians [11] use a video object tracker [7] to associate masks across different views. However, in scenarios with significant changes in camera poses between frames, such approaches struggle to maintain accuracy. OmniSeg3D [33] and Click-Gaussians [9] use contrastive learning-based method to attach each Gaussian with a feature, while the features of two Gaussians are encouraged to be similar if their projections are within the same segmentation mask. However, such methods do not provide segmentation mask labels for each segmented Gaussian group.
3 Proposed Method
Figure 3: Overview of Gaga. Gaga reconstructs 3D scenes using Gaussian Splatting and adopts any open-world model to generate 2D segmentation masks. To eliminate the 2D mask label inconsistency, we design a mask association process, where a 3D-aware memory bank is employed to assign a consistent group ID across different views to each 2D mask based on the 3D Gaussians projected to that mask (Sec. 3.2). Specifically, we find the corresponding Gaussians projected to 2D mask and assign the mask with the group ID in the memory bank with the maximum overlapped Gaussians (Eq. 5) After 3D-aware mask association process, we use masks with multi-view consistent group IDs as pseudo labels to train an identity encoding on each 3D Gaussian for segmentation rendering.
3.1 Preliminaries
Gaussian Splatting. Gaussian Splatting [13] has significantly advanced the 3D representation field by combining the benefits of implicit and explicit 3D representations. Specifically, a 3D scene is parameterized by a set of 3D Gaussians $\{G_{i}\}$. Each Gaussian $G_{i}=\{p_{i},s_{i},q_{i},\alpha_{i},c_{i}\}$ is defined by its position $p_{i}=\{x,y,z\}\in\mathbb{R}^{3}$, scale $s_{i}\in\mathbb{R}^{3}$, orientation $q\in\mathbb{R}^{4}$, opacity $\alpha\in\mathbb{R}$ and color features $c$ encoded by spherical harmonics (SH) coefficients.
Gaussian Splatting employs the splatting pipeline, wherein 3D Gaussians are projected onto the 2D image space using the world-to-frame transformation matrix corresponding to each camera pose. Gaussians projected to the same coordinates $(x,y)$ (represented as $i\in N$) are blended in depth order and weighted by their opacity $\alpha$ to produce the color $c_{x,y}$ of each pixel:
$c_{x,y}=\sum_{i\in N}c_{i}\alpha_{i}\prod_{j=1}^{i-1}(1-\alpha_{j}).$ (1)
Identity Encoding. Identity encoding ($e_{i}$[32] aims to assign a universal label to each 3D Gaussian for 3D scene segmentation. An identity encoding is a 16-dimensional feature vector attached to a 3D Gaussian. It can be decoded to a segmentation mask ID through a combination of linear and SoftMax layers. The segmentation label of a pixel (denoted as $m_{x,y}$) can then be computed by:
$m_{x,y}=\arg\max\{L(\sum_{i\in N}e_{i}\mathbf{\alpha}_{i}\prod_{j=1}^{i-1}(1-\mathbf{\alpha}_{j}))\}.$ (2)
The learning process for identity encoding involves optimizing the feature vectors by comparing the rendered masks with the ground truth 2D segmentation masks.
3.2 Grouping via 3D-aware Memory Bank
Given a set of posed images, we aim to reconstruct a 3D scene with semantic labels for segmentation rendering. To this end, we first leverage Gaussian Splatting for scene reconstruction. We then employ open-vocabulary 2D segmentation methods such as SAM [16] or EntitySeg [23] to predict class-agnostic segmentation for each input view.
However, because the segmentation model processes each input view independently, the resulting masks are not naturally multi-view consistent. To resolve this issue, existing works [32, 11] carry out a mask association process that tries to consolidate inconsistent segmentation masks from different views. For example, GaussianGrouping [32] and CoSSegGaussians [11] apply a video tracker to associate inconsistent 2D masks of different views, assuming similarity between nearby input views. This assumption, however, may not hold for all 3D scenes, particularly when the input views are sparse, as demonstrated in Fig. 2.
Gaga is inspired by the fundamental disparity between the task of mask association across multiple views and object tracking in a video: the latter does not exploit 3D information in the scene. To reliably generate consistent masks across different views, we propose a method that leverages 3D information without relying on any assumptions about the input views. Our key insight is that masks belonging to the same object in different views shall correspond to the same Gaussians in the 3D space. Consequently, these Gaussians should be grouped together and assigned with an identical group ID.
Corresponding Gaussians of Mask. Based on this intuition, we first associate each 2D segmentation mask with its corresponding 3D Gaussians. Specifically, we splat all 3D Gaussians onto the camera frame, using the camera pose of each input view. Subsequently, for each mask within the image, we identify 3D Gaussians whose centers are projected within that mask. Those Gaussians should be identified as representatives of the mask in 3D and can be used as guidance for associating masks from different views.
Notably, 2D segmentation masks typically describe the shape of foreground objects as observed from the current camera view. However, a significant proportion of Gaussians do not contribute to the pixels in the 2D segmentation mask, as they represent objects in the far back of the 3D scene. In Fig. 4, we show one example in column 3.
To address this challenge, we propose incorporating depth information as guidance to select Gaussians corresponding to foreground objects. We first render the depth map for each view. Given a mask $m$ in that view, we extract the corresponding depth values, denoted as $D$, and compute its minimum and maximum depths, $D_{\min}$ and $D_{\max}$, respectively. To mitigate the impact of inaccurate masks, we filter out outlier depths using the following procedure. We compute the first quartile ($Q_{1}$) and third quartile ($Q_{3}$) as:
$Q_{1}=\text{Quantile}(D,0.25),\quad Q_{3}=\text{Quantile}(D,0.75).$ (3)
The interquartile range (IQR) is then defined as $\text{IQR}=Q_{3}-Q_{1}$. To refine the depth range, we define the maximum inlier depth as $D_{\text{max\_inlier}}=Q_{3}+f\cdot\text{IQR}$, where $f$ is the outlier factor, set to 1.0 if not explicitly defined. The inliers’ depth range is then given by:
$S=\{z\in P\mid D_{\min}<z<D_{\text{max\_inlier}}\}.$ (4)
Finally, we select Gaussians that are projected within mask $m$ and lie inside the inlier depth range $S$ as the corresponding Gaussians for mask $m$. As shown in Fig. 4 (column 4), this approach produces a more accurate representation of the foreground objects, as it takes into account the spatial distribution of the Gaussians relative to the camera pose. By leveraging depth information, we can effectively filter out Gaussians that do not contribute to the foreground objects, thereby improving the accuracy of our 3D segmentation.
Figure 4: Corresponding Gaussians with Depth Guidance. In View 1, we select Gaussians (colored in red) that are splatted inside the mask of the bulldozer. As shown in column 3, some of these Gaussians not only belong to the bulldozer but also represent background objects, as seen in View 2. To refine the selection, we render the depth map and retain only the Gaussians within the specified depth region, ensuring they correspond to the bulldozer’s mask. As shown in column 4, the final selection with depth guidance consists primarily of Gaussians belonging to the bulldozer.
3D-aware Memory Bank. Next, to collect and categorize 3D Gaussians into groups and use them to associate masks across different views, we introduce a 3D-aware Memory Bank (see Fig. 3). Given a set of images, we initialize the 3D-aware Memory Bank by storing the corresponding Gaussians of each category in the first image into an individual group and labeling each mask with a group ID the same as its mask label. For each 2D mask of the subsequent view, we first determine its corresponding Gaussians as outlined above. We then either assign these Gaussians to an existing group within the memory bank or establish a new one if they do not share similarities with existing categories in the memory bank. In the following, we elaborate on the details of this assignment process.
Group ID Assignment via Gaussian Overlap. To assign each mask a group ID, we aim to find if the current mask has a significant amount of overlapped Gaussians with any groups in the memory bank. We define the similarity between two sets of 3D Gaussians based on their shared 3D Gaussian ratio. Specifically, given the 3D Gaussians corresponding to a 2D mask $m$ (denoted as $\mathcal{G}(m)$ as described above) and the Gaussians of category $i$ (denoted as $\mathcal{G}_{i}$) in the memory bank, we identify their shared Gaussians as $\mathcal{G}(m)\cap\mathcal{G}_{i}$ (i.e., Gaussians of the same indices), we then compute the overlap as the ratio of the number of shared Gaussians to number of all Gaussians within mask $m$:
$Overlap(m,i)=\frac{\#(\mathcal{G}(m)\cap\mathcal{G}_{i})}{\#\mathcal{G}(m)}.$ (5)
Note that we do not use the IoU of $\mathcal{G}(m)$ and $\mathcal{G}_{i}$ for the following reason: as more frames are processed, more Gaussians will be added to the memory bank, which means that $\mathcal{G}_{i}$ will become larger. Consequently, the IoU threshold needs to be adjusted frequently. In contrast, our overlapping formulation in Eq. 5 is independent of Gaussian number in the memory bank, avoiding manual threshold adjustment.
Suppose category $i$ has the highest overlap with mask $m$ among all categories in the memory bank, and this overlap value is above a threshold. In this case, we assign the group ID of mask $m$ as $i$ and add the non-overlapped Gaussians in the $i_{th}$ category by $\mathcal{G}_{i}=\mathcal{G}_{i}\cup\mathcal{G}(m)$. We establish a new group ID $j$ if none of the existing groups contains an overlap with mask $m$ above the overlapping threshold. We add $\mathcal{G}(m)$ into this new category in the Gaussian memory bank and assign mask $m$ with the new group ID $j$. Note that we ensure each Gaussian is added to only one group in the memory bank by tracking all the Gaussian indices already in the memory bank.
3.3 3D Segmentation and Applications.
After the group ID assigning process, masks projected by the same group of Gaussians are supposed to have the same group ID across different views. We splat the group IDs of 3D Gaussians in the memory bank to obtain pseudo ground truth segmentation masks, which are used to learn a 16-dimension feature vector (i.e., identity encoding [32]) for each Gaussian. Specifically, we initialize the identity encoding to all zeros for each 3D Gaussian. At each training iteration, we splat the identity encoding given a camera pose to formulate a 2D feature map. A linear layer followed by a SoftMax function are employed to predict a 2D segmentation map given the rendered feature map. Finally, we compute the Cross Entropy loss between the predicted segmentation maps and those obtained by rendering Gaussians in the memory bank, as discussed above.
After training, our segmentation-aware 3D Gaussians can be readily used for various downstream applications. For instance, we can render segmentation masks of novel views with consistent semantic labels for the same object across different camera poses. Furthermore, Gaussians can also be selected using their identity encoding and manipulated for 3D scene editing tasks, including removal, color-changing, position translation, etc., as will be demonstrated in Sec. 10 and supplementary material.
4 Experiments
4.1 Experimental Setup
Datasets. We experiment with various datasets across diverse scenarios to demonstrate the performance of Gaga. For quantitative comparison, we use a scene understanding dataset LERF-Mask [32], along with two indoor scene datasets: Replica [27] and ScanNet [10]. Additionally, we showcase the robustness of Gaga against variations in training image quantity by sparsely sampling the Replica dataset. We present visual comparison results on the commonly used scene reconstruction dataset, MipNeRF 360 [1]. The quantitative and qualitative evaluations are conducted on the test set, i.e., novel view synthesis results. Details about datasets can be found in the supplementary material.
Evaluation Metrics. Similarly to prior work [32], mIoU and boundary IoU (mBIoU) are used to evaluate the LERF-Mask dataset. To evaluate the multi-view consistency of 3D segmentation masks, we select eight scenes from the Replica dataset and seven scenes from the ScanNet dataset and use the ground-truth panoptic segmentation for evaluation, disregarding class information. The evaluation process is detailed in the supplementary material. Note that only Gaussian Grouping and Gaga provide 3D mask IDs to render multi-view consistent segmentation masks, we only compare these two methods. To handle differences between predicted and ground truth mask labels, we calculate the best linear assignment based on IoU. Moreover, with IoU = 0.5 as the criterion, we report precision and recall to further evaluate the accuracy of predicted masks. Please note that all numbers in the tables are expressed as percentages.
Implementation Details. We use SAM [16] and EntitySeg [23] with the Hornet-L backbone to obtain open-world 2D segmentation. We preprocess the generated raw masks following the method outlined in [23], prioritizing those with higher confidence scores by ranking them accordingly. Masks with confidence scores below 0.5 are discarded. For all experiments, we train vanilla Gaussian Splatting [13] for 30K iterations and train the identity encoding for 10K iterations with all other parameters frozen. We train baseline methods following their official guidance for 40K iterations for fair comparisons. We set the overlap threshold for declaring a new group ID as 0.1.
figurines ramen teatime
Model mIoU mBIoU mIoU mBIoU mIoU mBIoU
SA3D [3]* 24.9 23.8 7.4 7.0 42.5 39.2
LERF [14]* 33.5 30.6 28.3 14.7 49.7 42.6
LEGaussians [25] 34.6 32.6 31.4 18.8 42.8 35.3
LangSplat [24] 61.9 60.9 61.9 54.7 59.8 52.7
Gaussian Grouping [32]* 69.7 67.9 77.0 68.7 71.7 66.1
OmniSeg3D [33] 85.0 83.7 83.6 75.5 69.8 63.8
Gaga (Ours) 92.3 90.8 72.0 63.3 71.2 68.4
Table 1: Quantitative results for 3D query tasks on LERF-Mask. Gaga outperforms previous approaches, showcasing favorable performance in mIoU and BIoU with SAM. * denotes the results are reported in [32].
Figure 5: Qualitative results on LERF-Mask. Our rendered segmentation exhibits fewer artifacts and delivers more accurate segmentation results than both prior 3D class-agnostic segmentation works and language embedding works.
4.2 Open-vocabulary 3D Query on LERF-Mask
We compare our method with 6 state-of-the-art methods on 3D scene understanding: LERF [14], LEGaussians [25], LangSplat [24], SA3D [3], Gaussian Grouping [32] and OmniSeg3D [33]. The first three methods focus on CLIP feature embedding. We calculate the relevancy between rendered CLIP features and query text features following their official implementation. The last three methods leverage 2D segmentation masks predicted by SAM [16]. Since SAM does not support language prompts, SA3D, Gaussian Grouping and Gaga adopt Grounding DINO [19] to identify the mask ID in a 2D image and pick the corresponding 3D mask. OmniSeg3D does not provide mask IDs, so we use Grounding DINO to identify a pixel that semantically aligns with the text prompt and select the corresponding 3D mask based on a similarity threshold of 0.9.
Tab. 1 illustrates that Gaga achieves superior results in mIoU and mBIoU compared to previous methods utilizing SAM as 2D segmentation method. As shown in Fig. 5, the visualization results of 3D query tasks with prompts “paper napkin” and “bag of cookies” further demonstrate the advancement of Gaga, as Gaga provides more clear segmentation masks without artifacts.
2D Seg. Replica ScanNet
Model Method IoU Precision Recall IoU Precision Recall
GG [32] EntitySeg 35.90 14.07 31.57 39.54 6.88 36.56
Gaga (Ours) 41.45 59.74 45.59 43.08 35.46 49.73
GG [32] SAM 21.76 25.00 19.72 34.24 18.70 32.61
Gaga (Ours) 46.21 39.53 50.62 45.06 22.88 51.02
Table 2: Quantitative results on Replica and ScanNet. Gaga performs well with both 2D segmentation methods on two datasets. The performance of Gaussian Grouping (noted as GG) varies significantly with different 2D segmentation methods, whereas Gaga consistently delivers stable performance.
Figure 6: Qualitative results on Replica and ScanNet. Gaga consistently predicts more accurate segmentation masks. Gaussian Grouping (noted as GG) often covers the same object with different masks (row 1, 3 and 5), creating large empty regions (row 1, 2 and 3), or misidentifying similar instances (row 4 and 5).
2D Seg. Method EntitySeg SAM
Model Training Data IoU $\uparrow$ IoU Drop $\downarrow$ IoU $\uparrow$ IoU Drop $\downarrow$
GG [32] 30% 28.42 20.85 17.02 21.78
Gaga (Ours) 40.40 6.22 42.53 5.61
GG [32] 20% 24.56 31.35 16.02 26.38
Gaga (Ours) 39.00 9.46 42.14 6.47
GG [32] 10% 20.62 42.56 13.97 35.78
Gaga (Ours) 35.92 16.60 39.06 13.30
GG [32] 5% 10.00 72.15 6.77 68.87
Gaga (Ours) 29.67 31.11 31.87 29.27
Table 3: Quantitative results on Replica with limited data. Gaga consistently outperforms Gaussian Grouping with both 2D segmentation methods. The percentage of IoU drop shows that Gaga has greater robustness against reductions in training data.
4.3 3D Segmentation on Replica and ScanNet
Tab. 2 presents the quantitative comparison results on Replica and ScanNet datasets. Gaga exhibits better performance on both datasets regardless of the 2D segmentation model used, showcasing its stability across different datasets and models and consistently achieving high performance. Qualitative results are shown in Fig. 6. Row 1 and 2 depict the visualizations from the Replica dataset, while rows 3-5 showcase results from the ScanNet dataset. Gaussian Grouping [32] frequently assigns different mask IDs to the same object, resulting in inconsistent masks and empty regions. Row 4 shows that Gaussian Grouping struggles to distinguish similar objects, whereas Gaga accurately identifies each object by leveraging 3D information.
4.4 3D Segmentation with Limited Data on Replica
To demonstrate the robustness of Gaga against changes in training image quantity, we sparsely sample the Replica training set with ratios of 0.3, 0.2, 0.1, and 0.05. As depicted in Tab. 3, Gaga consistently exhibits superior performance in terms of IoU, with approximately a 15% advantage using EntitySeg and a 25% advantage using SAM. Remarkably, when utilizing SAM, Gaga surpasses fully trained Gaussian Grouping with just 5% of the training data by more than 10% (31.87% vs. 21.76%). We also compute the IoU drop compared to using all training images by:
$IoU\ Drop(x\%)=\frac{IoU(100\%)-IoU(x\%)}{IoU(100\%)},$ (6)
where $IoU(x\%)$ denotes the IoU achieved when $x\%$ of the training data is used. Compared to Gaussian Grouping, Gaga exhibits less sensitivity to decreases in the number of training images, as evidenced by smaller values in IoU drop. Visual results are shown in Fig. 7. With just 5% of the training data, Gaga delivers accurate segmentation masks, whereas Gaussian Grouping fails to provide masks for a significant portion of objects due to inaccurate tracking.
Figure 7: Qualitative results on Replica with limited data. The visualizations depict samples when using only 5% of the training data. Even with limited data, Gaga consistently produces high-quality segmentation. In contrast, Gaussian Grouping struggles to track objects accurately and leaves significant empty regions.
4.5 3D Segmentation on MipNeRF 360
Figure 8: Qualitative results on MipNeRF 360. Gaga provides superior masks with finer details (row 1 and 2), fewer artifacts and empty regions (row 1, 3 and 4), more consistent object masks across multi-views (wall in row 1, tablecloth in row 3).
We further showcase the performance of Gaga on the diverse MipNeRF 360 dataset. We provide visualization comparison on both indoor and outdoor scenes with Gaussian Grouping [32] in Fig. 8, using SAM [16] (row 1, 2) and EntitySeg [23] (row 3, 4). We display two images for each scene to assess the consistency across different views. Gaga offers more detailed segmentation, while segmentation masks generated by Gaussian Grouping exhibit severe artifacts and empty regions. Additionally, inconsistency across two views exists in the rendering results of Gaussian Grouping, as shown in row 1 and row 3.
4.6 Application: Scene Manipulation
Figure 9: Scene manipulation results on MipNeRF 360 and Replica. Gaga accurately identifies the cushion part of the footstool, whereas Gaussian Grouping colors it entirely. For object removal and translation tasks, Gaga generates more precise 3D entities with fewer artifacts, resulting in better visual performance.
Figure 10: Visual comparison of mask association methods. Gaga with 3D-aware memory bank achieves a superior visual quality and closer to the ground truth. Notice that the Video Tracker baseline mislabels the wall and floor, Our Memory Bank (w. All Gaussians) mislabels the floor.
Gaga achieves high-quality and multi-view consistent 3D segmentation, beneficial for tasks like scene manipulation. We demonstrate the effectiveness of this application on the MipNeRF 360 [1] and Replica [27] datasets (see Fig. 9). In the left image, we change the color of the footstool’s cushion and remove the stuffed animal on the armchair. Gaga accurately identifies 3D Gaussians representing the cushion, while Gaussian Grouping fails, coloring the entire footstool maroon along with part of the sofa and some floating Gaussians. Gaga also effectively groups and removes the entire part of the stuffed animal on the sofa with minimal artifacts, whereas Gaussian Grouping leaves many floating Gaussians. Similar results are observed in experiments involving the position shifting of a chair in the right image.
4.7 Ablation Study on Mask Association Method
Baseline IoU Precision Recall
SAM (Upper Bound) [16] 60.89 57.07 67.16
Linear Assignment [26] 1.92 1.54 0.58
w/o. Mask Association 8.81 3.19 2.16
Video Tracker [32] 21.76 25.00 19.72
Our Memory Bank (All Gaussians) 42.30 40.48 46.51
Our Memory Bank 46.21 39.53 50.62
Table 4: Ablation study on mask association methods. Our mask association method with the 3D-aware memory bank surpasses the previous video tracker baseline on all metrics.
We conduct ablation studies to evaluate the effectiveness of the proposed mask association method on the Replica dataset. The baselines for comparison include: (1) Linear Assignment: Cost-based linear assignment mask association proposed in Panoptic Lifting [26]. (2) w/o. Mask Association: Lifting inconsistent 2D masks to 3D. (3) Video Tracker[32] is employed as a representative method. (4) Our Memory Bank (w. All Gaussians): same as Gaga except that it selects all Gaussians splatted to the mask as its corresponding Gaussians. (5) Our Memory Bank: i.e., Gaga. We also add a baseline SAM (Upper Bound), which uses SAM to process rendered RGBs from Gaussian Splatting and evaluate on each single frame without considering multi-view consistency. This baseline can serve as an upper bound to show the inherent difference between class-agnostic and panoptic segmentation.
Quantitative results in Tab. 4 indicate that Gaga with the 3D-aware memory bank achieves superior performance compared to the previous method with video tracker. The cost-based linear assignment method used in [26] is not suitable for class-agnostic segmentation tasks, as it involves a significantly larger number of masks compared to panoptic segmentation. Comparison with Our Memory Bank (w. All Gaussians) baseline demonstrates the effectiveness of using depth maps to select corresponding 3D Gaussians for each mask. We also show the visual comparison in Fig. 10.
5 Conclusions
We introduce Gaga, a framework that reconstructs and segments open-world 3D scenes using inconsistent 2D masks predicted by zero-shot segmentation models. Gaga employs a 3D-aware memory bank to categorize 3D Gaussians and establishes mask association across different views by identifying the overlap between Gaussians projected to each mask. Results on various datasets show that Gaga outperforms previous methods with superior segmentation accuracy, multi-view consistency, and reduced artifacts. Additional application in scene manipulation further highlights Gaga’s high segmentation accuracy and practical utility.
Supplementary Material
1 Overview
In this supplementary document, we provide further experimental results, including more results on scene manipulation and sparse view setting in Sec. 3. We then delve into more experimental details of the datasets, metrics and implementation in Sec. 4. More ablation studies are shown in Sec. 5 and limitations are discussed in Sec. 6.
2 Supplementary Video
Please watch the supplementary demo video for a comprehensive introduction and visual comparison between our method Gaga and the current state-of-the-art methods. The video features additional qualitative comparisons and an animation illustration of Gaga.
Figure 11: Scene manipulation results on MipNeRF 360 and LERF-Mask. Gaga accurately identifies the flowerpot without affecting the color of the plant. Notice that Gaussian Grouping [32] creates a cyan region on the wooden door behind. For the object removal and duplication tasks, Gaga can also provide more accurate results with fewer artifacts.
3 Supplementary Experimental Results
3.1 Results on Scene Manipulation
Gaga can accurately segment the Gaussians of a 3D object and edit their properties. Using a pre-trained 3D Gaussian model with identity encoding, we employ the classifier trained with identity encoding to predict mask labels for each 3D Gaussian. Subsequently, we select 3D Gaussians sharing the same mask label as the target object and edit their properties for tasks like object coloring, removal, and position translation.
We provide additional results for the downstream scene manipulation task to further demonstrate the prospect of applying Gaga to real-world scenarios. On the “counter” scene of the MipNeRF 360 dataset [1], we change the color of the flowerpot to cyan and duplicate the glass jar. Gaussian Grouping [32] can not differentiate the plant and flowerpot, whereas Gaga generates a more accurate segmentation mask. Additionally, Gaga produces a clearer boundary and avoids artifacts on the iron tray when duplicating the glass jar.
In the “figurines” scene of the LERF-Mask dataset [32], we transform the yellow duck to blue and remove the red toy chair. Gaga precisely changes only the duck’s color without affecting other objects, and achieves a more thorough removal of the red toy chair.
3.2 Results on Sparsely Sampled Replica
Figure 12: Qualitative results on the sparsely sampled Replica. We showcase the novel view synthesis segmentation rendering results provided by Gaussian Grouping and Gaga as the percentage of training images employed decreases from 100% to 5%. Gaussian Grouping cannot correctly track the sofa under sparse views and fails to differentiate ceiling and wall, whereas Gaga consistently provides high-quality segmentation results.
We provide additional qualitative results for the experiment on the sparsely sampled replica dataset in Fig. 12. As the number of training images decreases, Gaussian Grouping produces more empty regions, e.g. the sofa, due to difficulties in accurate tracking under sparse views. Whereas Gaga exhibits a more robust performance against reductions in the number of images.
4 Experimental Details
4.1 Details on Datasets
We employ the official script from Gaussian Splatting [13] for colmap to acquire camera poses and the initial point cloud. Consequently, the actual number of images utilized in the experiment might be lower than expected due to colmap process failures. Please refer to Tab. 5 for the scene names used in the Replica and ScanNet datasets.
LERF-Mask Dataset [32]. LERF-Mask is based on the LERF dataset [14] and annotated with tasks and ground truth by the author of [32]. It contains 3 scenes: figurines, ramen, and teatime. For each scene, 6-10 objects are selected as text queries, and Grounding DINO [19] is utilized to select the mask ID from the rendered segmentation.
Replica Dataset [27]. We select 8 scenes from the entire Replica Dataset the same as [35]. We use the rendered results provided by authors of [35] and follow their data processing process: for each scene, we uniformly select 20% images as training data and 20% images as test data from all rendered RGB images. This results in 180 training images and 180 test images for each scene.
Sparsely Sampled Replica Dataset. For the same 8 scenes as the previous experiment, we randomly sample 30%, 20%, 10%, and 5% of the total 180 training images, resulting in 54, 36, 18, and 9 training images for each task, respectively. The number of test images remains at 180.
ScanNet Dataset [10]. DM-NeRF [29] selects 8 scenes from the entire ScanNet dataset. Each scene has approximately 300 images for training and about 100 images for testing. We utilize 7 out of the 8 scenes, excluding ”scene 0024_00” due to the subpar 3D reconstruction results in both Gaussian Splatting [13] and Gaussian Grouping [32].
MipNeRF 360 Dataset [1]. We downsample the images by a factor of 4, consistent with the setting in [32], to accommodate the large size of the original images. For novel view synthesis evaluation, we set the sample step at 8, the same as the setting in [13].
Dataset Scene Name
Replica office 0 office 1 office 2 office 3
office 4 room 0 room 1 room 2
ScanNet scene 0010_00 scene 0012_00 scene 0033_00 scene 0038_00
scene 0088_00 scene 0113_00 scene 0192_00
Table 5: Selected scenes in Replica and ScanNet. We select 8 scenes from the Replica dataset following [35], and 7 scenes from the ScanNet dataset following [29].
4.2 Details on Evaluation Metrics
Given the disparate mask label assignments between the ground truth segmentation and the predicted segmentation for 3D objects, we find the best linear assignment between the labels based on IoU for quantitative evaluation. Subsequently, we employ IoU $>$ 0.5 as the criterion for precision and recall calculations. We outline the pseudocode for the evaluation procedure in Algorithm 1. Note that all annotated segmentation masks are unavailable during training and are only accessible during evaluation as ground truth.
Algorithm 1 Evaluation Metrics
Input pred_masks and gt_masks are represented in binary format with shape ($n_{image}$, $n_{mask}$, $h$, $w$), where $n_{image}$ is the number of test images, $n_{mask}$ is the number of predicted or ground truth masks, $h$, $w$ are the height and width of test images.
We use scipy.optimize.linear_sum_assignment to solve the linear assignment problem.
   Function evaluate(pred_masks, gt_masks)
    Input: pred_masks (torch.bool), gt_masks (torch.bool)
    Output: iou (torch.float), precision (torch.float), recall (torch.float)
   assert len(gt_masks) == len(pred_masks)
   $n_{image}$leftarrow len(gt_masks)
   $n_{pred}$leftarrow pred_masks.shape[1]
   $n_{gt}$leftarrow gt_masks.shape[1]
   iou_matrix $\leftarrow$ torch.zeros(($n_{gt}$, max($n_{gt}$, $n_{pred}$)))
   for $i$ in $n_{gt}$ do
      for $j$ in $n_{pred}$ do
         iou_list $\leftarrow$ []
         for $k$ in $n_{image}$ do
            iou_list.append(IoU(gt_masks[$k$][$i$], pred_masks[$k$][$j$]))
         end for
         iou_matrix[$i$][$j$] $\leftarrow$ iou_list.mean()
      end for
   end for
   $gt\_indices$, $pred\_indices$leftarrow linear_assignment(iou_matrix)
   paired_iou $\leftarrow$ iou_matrix[$gt\_indices$][$pred\_indices$]
   iou $\leftarrow$ paired_iou.mean()
   $n_{correct}$leftarrow torch.sum(paired_iou $>$ 0.5)
   precision $\leftarrow$ $\frac{n_{correct}}{n_{pred}}$
   recall $\leftarrow$ $\frac{n_{correct}}{n_{gt}}$
   return iou, precision, recall
4.3 Further Implementation Details
For training vanilla 3D Gaussians, we maintain the same parameter setting as [13]. To train the identity encoding, we freeze all the other attributes of Gaussians and use the same parameter setting as [32]. The identity encoding has 16 dimensions, and the rendered 2D identity encoding is in the shape of 16 $\times$ $h$$\times$ $w$, where $h$ and $w$ represent the height and width of the image. The classifier for predicting mask ID given the 2D identity encoding and selecting Gaussians for editing given the 3D identity encoding shares the same architecture, with 16 input channels. The number of output channels equals the number of groups in the 3D-aware memory bank after associating all images. All experiments are conducted on a single NVIDIA RTX 6000 Ada GPU.
figurines ramen teatime
Overlap Threshold mIoU mBIoU mIoU mBIoU mIoU mBIoU
0.01 79.6 77.5 72.0 63.1 71.2 68.2
0.05 91.4 89.8 72.3 63.3 70.2 67.5
0.1 92.3 90.1 72.0 63.3 71.2 68.4
0.2 85.8 84.3 51.3 48.7 71.7 69.4
0.3 85.7 84.0 45.6 43.3 71.6 69.0
Table 6: Ablation study on the overlap threshold. If the overlap between the current mask and all groups in the memory bank falls below this threshold, we add this mask to the memory bank as a new group. Results indicate that the default setting of 0.1 generally yields better outcomes.
5 Supplementary Ablation Study
5.1 Ablation Study on Overlap Threshold
We conduct additional ablation studies on the Gaussian overlap threshold using the Replica dataset [27], with SAM [16] as the 2D segmentation model. During the group ID assignment process, if none of the existing groups in the memory bank exceed the overlap threshold with the current mask, we add the mask as a new group, indicating the discovery of a new 3D object.
Tab. 6 presents our ablation study on the overlap threshold. When set to 0.01, new groups are rarely created, favoring association with existing ones. This yields the highest precision but results in lower IoU performance. Conversely, a threshold of 0.3 leads to frequent creation of new group IDs, achieving the best IoU but significantly reducing precision. To balance performance across all three metrics, we set the threshold to 0.1.
5.2 Comparison on Mask Association Methods
Figure 13: Visual comparison between different mask association methods. Our 3D-aware memory bank offers more detailed associated masks, accurately tracks identical objects in the scene and assigns them different mask IDs. Conversely, video object tracker leaves empty regions in positions where it cannot track masks, and it struggles to provide consistent masks for the same object across views.
We present visual comparison results for two mask association methods, video tracker [7] utilized by [32] and Gaga’s 3D-aware memory bank, in Fig. 13. In the “garden” scene of the MipNeRF 360 dataset, the video tracker struggles to track objects in the background, whereas Gaga provides associated results for each mask. For the scene in the ScanNet dataset, the video tracker fails to distinguish between four identical sofas, resulting in multiple masks for the same object. Additionally, it assigns different mask IDs to the table in two views. In contrast, Gaga precisely locates each object, leading to improved mask association results and better pseudo labels for training segmentation features.
6 Limitations
Though Gaga achieves SOTA performance compared to existing works, there are a few limitations and future works. First, the optimization process of identity encoding and the rest of the Gaussian parameters are independent, this is because we need to first train 3D Gaussians to acquire their spatial location for mask association. While this pipeline allows for the utilization of any pre-trained 3D Gaussians as input without the need to re-train the entire scene, it does require additional training steps. We aim to enable the joint processing of mask association and identity encoding training in future works.
Secondly, artifacts may occur in the segmentation rendered by Gaga due to inherent inconsistency in the 2D segmentation. For example, an object might be depicted as one mask in the initial view but as two separate masks in subsequent views. This ambiguity introduces challenges to our mask association process. Preprocessing steps such as dividing, merging, or reshaping the 2D segmentation masks could potentially resolve this issue and improve grouping results.

### v3

@@
-Gaga: Group Any Gaussians via 3D-aware Memory Bank
+Group Any Gaussians via 3D-aware Memory Bank
 Abstract
@@
 Gaga produces accurate 3D object segmentation, achieving high-quality results for downstream applications such as scene manipulation (e.g.
-changing the cushion’s color of the footstool to maroon).
+changing the cushion’s color of the footstool to maroon)
 1 Introduction
+Effective open-world 3D segmentation is essential for scene understanding and manipulation.
+Despite notable advancements in 2D open-world segmentation techniques, exemplified by Segment Anything (SAM) sam and EntitySeg entityseg , extending these methodologies to the realm of 3D encounters the challenge of ensuring consistent mask label assignment across multi-view images.
+Specifically, masks of the same object across different views may have different mask label IDs, as the 2D class-agnostic segmentation model independently processes the multi-view images.
+Naively lifting these inconsistent masks to 3D introduces ambiguity and leads to inferior results in 3D scene segmentation.
 Figure 2: Comparison of rendered segmentation.
-Contrastive learning-based methods, such as OmniSeg [33], do not provide unique mask labels to each segmentation group, leading to inconsistencies across multiple views (e.g., the coffee table).
-Gaussian Grouping [32] addresses multi-view segmentation by utilizing a video tracker, but it often misidentifies objects when similar items are present (e.g., the leather sofa) and struggles with significant camera perspective changes.
+Contrastive learning-based methods, such as OmniSeg omniseg3d , do not provide unique mask labels to each segmentation group, leading to inconsistencies across multiple views (e.g., the coffee table).
+Gaussian Grouping gg addresses multi-view segmentation by utilizing a video tracker, but it often misidentifies objects when similar items are present (e.g., the leather sofa) and struggles with significant camera perspective changes.
 In contrast, Gaga ensures multi-view consistent segmentation masks, overcoming these limitations.
-Effective open-world 3D segmentation is essential for scene understanding and manipulation.
-Despite notable advancements in 2D open-world segmentation techniques, exemplified by Segment Anything (SAM) [16] and EntitySeg [23], extending these methodologies to the realm of 3D encounters the challenge of ensuring consistent mask label assignment across multi-view images.
-Specifically, masks of the same object across different views may have different mask label IDs, as the 2D class-agnostic segmentation model independently processes the multi-view images.
-Naively lifting these inconsistent 2D masks to 3D introduces ambiguity and leads to inferior results in 3D scene segmentation.
-Existing works [33, 15, 32, 11, 9] address this issue either by contrastive learning or video object tracking.
-One line of works build upon 3D scenes represented as 3D Gaussians [13].
+Existing works omniseg3d ; garfield ; gg ; cosseg ; clickgaussian address this issue either by contrastive learning or video object tracking.
+One line of work builds upon 3D scenes represented as 3D Gaussians 3dgs .
 In this approach, a feature vector is learned for each 3D Gaussian through a contrastive learning paradigm.
 The method encourages similar features for 3D Gaussians whose projections fall within the same segmentation mask, while pushing apart those belonging to different segmentation regions.
-Segmentation maps for each view are then obtained by clustering these feature vectors.
+The segmentation maps for each view are then obtained by clustering these feature vectors.
 However, these methods do not assign a specific mask label to each 3D segmentation group, resulting in inconsistent multi-view segmentation, as shown in Fig.
 2.
-Alternatively, some methods [32, 11] tackle multi-view inconsistency in 3D scene segmentation by treating multi-view images as a video sequence and adopting an off-the-shelf video object tracking method [7].
-Nevertheless, this design assumes minimal view changes between multi-view images—a condition that may not always hold in real-world 3D scenes, particularly when input views are sparse.
+Alternatively, some methods gg ; cosseg tackle multi-view inconsistency in 3D scene segmentation by treating multi-view images as a video sequence and adopting an off-the-shelf video object tracking method deva .
+Nevertheless, this design assumes minimal view changes between multiview images, a condition that may not always hold in real-world 3D scenes, particularly when input views are sparse.
 Consequently, these methods often struggle with similar objects or occluded objects that intermittently disappear and reappear in the sequence, as shown in Fig.
 2.
-In this paper, we identify a fundamental limitation in existing open-world 3D scene segmentation methods that leads to 3D inconsistency: the inadequate exploitation of 3D information inherently provided by the scene.
+We identify a fundamental limitation in existing open-world 3D scene segmentation methods that leads to 3D inconsistency: the inadequate exploitation of 3D information inherently provided by the scene.
 Our main intuition is that masks of the same object across different views shall correspond to the same group of 3D Gaussians.
@@
 Based on this intuition, we propose Gaga, which groups any 3D Gaussians and renders consistent 3D class-agnostic segmentation across different views.
-Given a collection of posed RGB images, we first employ Gaussian Splatting [13] to reconstruct the 3D scene and extract 2D masks using an open-world segmentation method [16, 23].
+Given a collection of posed RGB images, we first employ Gaussian Splatting 3dgs to reconstruct the 3D scene and extract 2D masks using an open-world segmentation method sam ; entityseg .
 Subsequently, we iteratively build a 3D-aware memory bank that collects and stores the Gaussians grouped by category.
-For each input view, we project each 2D mask into 3D space using camera parameters and search the memory bank for the category with the largest overlapping with the unprojected mask.
-Based on the degree of overlapping, we either assign the mask to an existing category or create a new one.
+For each input view, we project each 2D mask into a 3D space using camera parameters and search the memory bank for the category with the largest overlap with the unprojected mask.
+Based on the degree of overlap, we assign the mask to an existing category or create a new one.
 Finally, following the mask association process described above, we can get a set of multi-view consistent 2D segmentation masks.
-Finally, we learn a feature vector (i.e., identity encoding [32]) for each 3D Gaussian that encodes its category information.
-Specifically, we splat the feature vectors onto 2D image and decode a segmentation map through a linear layer.
+Finally, we learn a feature vector (i.e., identity encoding gg ) for each 3D Gaussian that encodes its category information.
+Specifically, we splat the feature vectors onto the 2D image and decode a segmentation map through a linear layer.
 The predicted masks are then compared with the segmentation masks obtained from our 3D-aware memory bank for supervision.
 This approach ensures that our identity encoding remains multi-view consistent, thanks to the consistent segmentation masks across different viewpoints.
-Our contributions include:
-  • We propose a framework that reconstructs and segments 3D scenes using inconsistent 2D masks generated by open-world segmentation models.
+The contributions of this work are: 1).
+We propose a framework for reconstructing and segmenting 3D scenes from inconsistent 2D masks generated by open-world segmentation models.
+2).
 To resolve the inconsistency of 2D masks across views, we design a 3D-aware memory bank that collects Gaussians of the same semantic group.
-This memory bank is then employed to align 2D masks across diverse views.
-  • We show that the proposed method can effectively leverage any 2D class-agnostic segmentation masks, making it readily applicable for synthesizing novel view images and segmentation masks.
-  • Comprehensive experiments on diverse datasets and challenging scenarios, including sparse input views, demonstrate the effectiveness of the proposed method both qualitatively and quantitatively.
+This memory bank is then used to align 2D masks across diverse views.
+3).
+We show that the proposed method can effectively leverage any 2D class-agnostic segmentation mask, making it readily applicable for synthesizing novel view images and segmentation masks.
+4.) Experiments on diverse datasets and challenging scenarios, including sparse input views, demonstrate the effectiveness of the proposed method qualitatively and quantitatively.
 2 Related Work
 Segment and Tracking Anything in 2D.
-Segment Anything (SAM) [16] and EntitySeg [23] demonstrate the effectiveness of large-scale training in image segmentation, thus establishing a pivotal foundation for open-world segmentation methods.
-Subsequent studies [31, 8, 7] further extend the applicability of SAM to video data by leveraging video object segmentation algorithms to propagate SAM masks.
-Conversely, acquiring data for training their 3D counterparts poses a challenge, given that existing large-scale 3D datasets with annotated segmentation [27, 10] primarily focus on indoor scenarios.
+Segment Anything (SAM) sam and EntitySeg entityseg demonstrate the effectiveness of large-scale training in image segmentation, thus establishing a pivotal foundation for open-world segmentation methods.
+Subsequent studies track ; sata ; deva further extend the applicability of SAM to video data by leveraging video object segmentation algorithms to propagate SAM masks.
+Conversely, acquiring data for training their 3D counterparts poses a challenge, given that existing large-scale 3D datasets with annotated segmentation replica ; scannet focus on indoor scenarios.
 NeRF-based 3D Segmentation.
-Neural Radiance Fields (NeRFs) [21] model scenes as continuous volumetric functions, learned through neural networks that map 3D coordinates to scene radiance.
+Neural Radiance Fields (NeRFs) nerf model scenes as continuous volumetric functions, learned through neural networks that map 3D coordinates to scene radiance.
 This approach facilitates the capture of intricate geometric details and the generation of photorealistic renderings, offering novel view synthesis capabilities.
-Semantic-NeRF [35] initiates the incorporation of semantic information into NeRFs and enables the generation of semantic masks for novel views.
+Semantic-NeRF semantic_nerf initiates the incorporation of semantic information into NeRFs and enables the generation of semantic masks for novel views.
 Note that semantic segmentation masks do not face the challenge of ambiguous mask ID across views.
-Numerous methods expand the scope by introducing instance modeling and matching instance masks relying on existing 3D bounding boxes [20, 12], resorting to the cost-based linear assignment during training [26, 29] or directly training instance-specific MLPs [17].
+Numerous methods expand the scope by introducing instance modeling and matching instance masks relying on existing 3D bounding boxes instance_nerf ; panoptic_nerf , resorting to the cost-based linear assignment during training panoptic_lifiting ; dm_nerf or directly training instance-specific MLPs pnf .
 However, most of these methods are developed based on ground truth segmentation and tailored for scene modeling within specific domains.
 They often entail high computational costs and lack substantial evidence of their performance in open-world scenarios.
-Leveraging SAM’s open-world segmentation capability, SA3D [3] endeavors to recover a 3D consistent mask by tracing 2D masks across adjacent views with user guidance.
+Leveraging SAM’s open-world segmentation capability, SA3D sa3d endeavors to recover a 3D consistent mask by tracing 2D masks across adjacent views with user guidance.
 Similarly, Chen et al.
-[6] distill SAM encoder features into 3D and query the decoder.
+interactive distill SAM encoder features into 3D and query the decoder.
 In contrast, Gaga achieves multi-view consistency without user intervention, offering segmentation for all objects rather than an instance.
-Garfield [15] densely samples SAM masks and trains a scale-conditioned affinity field supervised on the scale of each mask deprojected to 3D.
-However, this work focuses on 3D clustering rather than segmentation.
+Garfield garfield densely samples SAM masks and trains a scale-conditioned affinity field supervised on the scale of each mask deprojected to 3D, which focuses on 3D clustering rather than segmentation.
 Gaussian-based 3D Segmentation.
-As an alternative to NeRF and its variants [21, 4, 22, 28], Gaussian Splatting [13, 5, 30, 34] has recently emerged as a powerful approach to reconstruct 3D scenes via real-time radiance field rendering.
+As an alternative to NeRF and its variants nerf ; tensorf ; instant ; nerfstudio , Gaussian Splatting 3dgs ; chen2024survey ; recent ; Yu2023MipSplatting has recently emerged as a powerful approach to reconstruct 3D scenes via real-time radiance field rendering.
 By representing the scene as 3D Gaussians from posed images, it achieves photorealistic novel view synthesis with high reconstruction quality and efficiency.
 Additionally, manipulating 3D Gaussians for scene editing is more straightforward compared to NeRF’s representation.
-SAGA [2] renders a 2D SAM feature map and uses a SAM guidance loss to learn 3D segmentation from the ambiguous 2D masks.
-Similar to [3], this method requires user input and only provides segmentation for one object at a time.
-Feature 3DGS [36] distills LSeg [18] and SAM features to 3D Gaussians and decodes rendered features to obtain segmentation.
+SAGA saga renders a 2D SAM feature map and uses a SAM guidance loss to learn 3D segmentation from the ambiguous 2D masks.
+Similar to sa3d , this method requires user input and only provides segmentation for one object at a time.
+Feature 3DGS feature3dgs distills LSeg lseg and SAM features to 3D Gaussians and decodes rendered features to obtain segmentation.
 However, it fails to provide consistent segmentation across views.
-Gaussian Grouping [32] and CoSSegGaussians [11] use a video object tracker [7] to associate masks across different views.
+Gaussian Grouping gg and CoSSegGaussians cosseg use a video object tracker deva to associate masks across different views.
 However, in scenarios with significant changes in camera poses between frames, such approaches struggle to maintain accuracy.
-OmniSeg3D [33] and Click-Gaussians [9] use contrastive learning-based method to attach each Gaussian with a feature, while the features of two Gaussians are encouraged to be similar if their projections are within the same segmentation mask.
+OmniSeg3D omniseg3d and Click-Gaussians clickgaussian use contrastive learning-based method to attach each Gaussian with a feature, while the features of two Gaussians are encouraged to be similar if their projections are within the same segmentation mask.
 However, such methods do not provide segmentation mask labels for each segmented Gaussian group.
@@
 Specifically, we find the corresponding Gaussians projected to 2D mask and assign the mask with the group ID in the memory bank with the maximum overlapped Gaussians (Eq.
-5) After 3D-aware mask association process, we use masks with multi-view consistent group IDs as pseudo labels to train an identity encoding on each 3D Gaussian for segmentation rendering.
+3) After 3D-aware mask association process, we use masks with multi-view consistent group IDs as pseudo labels to train an identity encoding on each 3D Gaussian for segmentation rendering.
 3.1 Preliminaries
 Gaussian Splatting.
-Gaussian Splatting [13] has significantly advanced the 3D representation field by combining the benefits of implicit and explicit 3D representations.
+Gaussian Splatting 3dgs has significantly advanced the 3D representation field by combining the benefits of implicit and explicit 3D representations.
 Specifically, a 3D scene is parameterized by a set of 3D Gaussians $\{G_{i}\}$.
@@
 Gaussian Splatting employs the splatting pipeline, wherein 3D Gaussians are projected onto the 2D image space using the world-to-frame transformation matrix corresponding to each camera pose.
-Gaussians projected to the same coordinates $(x,y)$ (represented as $i\in N$) are blended in depth order and weighted by their opacity $\alpha$ to produce the color $c_{x,y}$ of each pixel:
-$c_{x,y}=\sum_{i\in N}c_{i}\alpha_{i}\prod_{j=1}^{i-1}(1-\alpha_{j}).$ (1)
-Identity Encoding.
-Identity encoding ($e_{i}$[32] aims to assign a universal label to each 3D Gaussian for 3D scene segmentation.
-An identity encoding is a 16-dimensional feature vector attached to a 3D Gaussian.
-It can be decoded to a segmentation mask ID through a combination of linear and SoftMax layers.
-The segmentation label of a pixel (denoted as $m_{x,y}$) can then be computed by:
-$m_{x,y}=\arg\max\{L(\sum_{i\in N}e_{i}\mathbf{\alpha}_{i}\prod_{j=1}^{i-1}(1-\mathbf{\alpha}_{j}))\}.$ (2)
-The learning process for identity encoding involves optimizing the feature vectors by comparing the rendered masks with the ground truth 2D segmentation masks.
-3.2 Grouping via 3D-aware Memory Bank
+Gaussians projected to the same coordinates $(x,y)$ (represented as $i\in N$) are blended in depth order and weighted by their opacity $\alpha$ to produce the color $c_{x,y}$ of each pixel: $c_{x,y}=\sum_{i\in N}c_{i}\alpha_{i}\prod_{j=1}^{i-1}(1-\alpha_{j})$.
+3.2 Grouping Any Gaussians via 3D-aware Memory Bank
 Given a set of posed images, we aim to reconstruct a 3D scene with semantic labels for segmentation rendering.
 To this end, we first leverage Gaussian Splatting for scene reconstruction.
-We then employ open-vocabulary 2D segmentation methods such as SAM [16] or EntitySeg [23] to predict class-agnostic segmentation for each input view.
+We then employ open-vocabulary 2D segmentation methods such as SAM sam or EntitySeg entityseg to predict class-agnostic segmentation for each input view.
 However, because the segmentation model processes each input view independently, the resulting masks are not naturally multi-view consistent.
-To resolve this issue, existing works [32, 11] carry out a mask association process that tries to consolidate inconsistent segmentation masks from different views.
-For example, GaussianGrouping [32] and CoSSegGaussians [11] apply a video tracker to associate inconsistent 2D masks of different views, assuming similarity between nearby input views.
+To resolve this issue, existing works gg ; cosseg carry out a mask association process that tries to consolidate inconsistent segmentation masks from different views.
+For example, GaussianGrouping gg and CoSSegGaussians cosseg apply a video tracker to associate inconsistent 2D masks of different views, assuming similarity between nearby input views.
 This assumption, however, may not hold for all 3D scenes, particularly when the input views are sparse, as demonstrated in Fig.
@@
 Our key insight is that masks belonging to the same object in different views shall correspond to the same Gaussians in the 3D space.
-Consequently, these Gaussians should be grouped together and assigned with an identical group ID.
+Consequently, these Gaussians should be grouped together and assigned an identical group ID.
 Corresponding Gaussians of Mask.
@@
 We compute the first quartile ($Q_{1}$) and third quartile ($Q_{3}$) as:
-$Q_{1}=\text{Quantile}(D,0.25),\quad Q_{3}=\text{Quantile}(D,0.75).$ (3)
+$Q_{1}=\text{Quantile}(D,0.25),\quad Q_{3}=\text{Quantile}.(D,0.75).$ (1)
+Figure 4: Corresponding Gaussians with Depth Guidance.
+In View 1, we select Gaussians (colored in red) that are splatted inside the mask of the bulldozer.
+As shown in column 3, some of these Gaussians not only belong to the bulldozer but also represent background objects, as seen in View 2.
+To refine the selection, we render the depth map and retain only the Gaussians within the specified depth region, ensuring they correspond to the bulldozer’s mask.
+As shown in column 4, the final selection with depth guidance consists primarily of Gaussians belonging to the bulldozer.
 The interquartile range (IQR) is then defined as $\text{IQR}=Q_{3}-Q_{1}$.
@@
 By leveraging depth information, we can effectively filter out Gaussians that do not contribute to the foreground objects, thereby improving the accuracy of our 3D segmentation.
-Figure 4: Corresponding Gaussians with Depth Guidance.
-In View 1, we select Gaussians (colored in red) that are splatted inside the mask of the bulldozer.
-As shown in column 3, some of these Gaussians not only belong to the bulldozer but also represent background objects, as seen in View 2.
-To refine the selection, we render the depth map and retain only the Gaussians within the specified depth region, ensuring they correspond to the bulldozer’s mask.
-As shown in column 4, the final selection with depth guidance consists primarily of Gaussians belonging to the bulldozer.
 3D-aware Memory Bank.
@@
 3 is independent of Gaussian number in the memory bank, avoiding manual threshold adjustment.
+In practice, We set the overlap threshold for declaring a new group ID as 0.1.
 Suppose category $i$ has the highest overlap with mask $m$ among all categories in the memory bank, and this overlap value is above a threshold.
@@
 After the group ID assigning process, masks projected by the same group of Gaussians are supposed to have the same group ID across different views.
-We splat the group IDs of 3D Gaussians in the memory bank to obtain pseudo ground truth segmentation masks, which are used to learn a 16-dimension feature vector (i.e., identity encoding [32]) for each Gaussian.
+We splat the group IDs of 3D Gaussians in the memory bank to obtain pseudo ground truth segmentation masks, which are used to learn a 16-dimensional feature vector (i.e., identity encoding gg ) for each Gaussian.
+Identity encoding ($e_{i}$gg aims to assign a universal label to each 3D Gaussian for 3D scene segmentation.
+It can be decoded to a segmentation mask ID through a combination of linear and SoftMax layers.
+The segmentation label of a pixel (denoted as $m_{x,y}$) is computed by:
+$m_{x,y}=\arg\max\{L(\sum_{i\in N}e_{i}\mathbf{\alpha}_{i}\prod_{j=1}^{i-1}(1-\mathbf{\alpha}_{j}))\}.$ (4)
+The learning process for identity encoding involves optimizing the feature vectors by comparing the rendered masks with the ground truth 2D segmentation masks.
 Specifically, we initialize the identity encoding to all zeros for each 3D Gaussian.
@@
 Datasets.
-We experiment with various datasets across diverse scenarios to demonstrate the performance of Gaga.
-For quantitative comparison, we use a scene understanding dataset LERF-Mask [32], along with two indoor scene datasets: Replica [27] and ScanNet [10].
+For quantitative comparison, we use a scene understanding dataset LERF-Mask gg , along with two indoor scene datasets: Replica replica and ScanNet scannet .
 Additionally, we showcase the robustness of Gaga against variations in training image quantity by sparsely sampling the Replica dataset.
-We present visual comparison results on the commonly used scene reconstruction dataset, MipNeRF 360 [1].
+We present visual comparison results on the commonly used scene reconstruction dataset, MipNeRF 360 mipnerf .
 The quantitative and qualitative evaluations are conducted on the test set, i.e., novel view synthesis results.
@@
 Evaluation Metrics.
-Similarly to prior work [32], mIoU and boundary IoU (mBIoU) are used to evaluate the LERF-Mask dataset.
+Similarly to prior work gg , mIoU and boundary IoU (mBIoU) are used to evaluate the LERF-Mask dataset.
 To evaluate the multi-view consistency of 3D segmentation masks, we select eight scenes from the Replica dataset and seven scenes from the ScanNet dataset and use the ground-truth panoptic segmentation for evaluation, disregarding class information.
@@
 Implementation Details.
-We use SAM [16] and EntitySeg [23] with the Hornet-L backbone to obtain open-world 2D segmentation.
-We preprocess the generated raw masks following the method outlined in [23], prioritizing those with higher confidence scores by ranking them accordingly.
+We use SAM sam and EntitySeg entityseg with the Hornet-L backbone to obtain open-world 2D segmentation.
+We preprocess the generated raw masks following the method outlined in entityseg , prioritizing those with higher confidence scores by ranking them accordingly.
 Masks with confidence scores below 0.5 are discarded.
-For all experiments, we train vanilla Gaussian Splatting [13] for 30K iterations and train the identity encoding for 10K iterations with all other parameters frozen.
+For all experiments, we train vanilla Gaussian Splatting 3dgs for 30K iterations and train the identity encoding for 10K iterations with all other parameters frozen.
 We train baseline methods following their official guidance for 40K iterations for fair comparisons.
-We set the overlap threshold for declaring a new group ID as 0.1.
+4.2 Open-vocabulary 3D Query on LERF-Mask
+Table 1: 3D query on LERF-Mask. Gaga outperforms previous approaches, showcasing favorable performance in mIoU and BIoU with both segmentation models.
 figurines ramen teatime
 Model mIoU mBIoU mIoU mBIoU mIoU mBIoU
-SA3D [3]* 24.9 23.8 7.4 7.0 42.5 39.2
-LERF [14]* 33.5 30.6 28.3 14.7 49.7 42.6
-LEGaussians [25] 34.6 32.6 31.4 18.8 42.8 35.3
-LangSplat [24] 61.9 60.9 61.9 54.7 59.8 52.7
-Gaussian Grouping [32]* 69.7 67.9 77.0 68.7 71.7 66.1
-OmniSeg3D [33] 85.0 83.7 83.6 75.5 69.8 63.8
+SA3D sa3d * 24.9 23.8 7.4 7.0 42.5 39.2
+LERF lerf * 33.5 30.6 28.3 14.7 49.7 42.6
+LEGaussians legaussians 34.6 32.6 31.4 18.8 42.8 35.3
+LangSplat langsplat 61.9 60.9 61.9 54.7 59.8 52.7
+Gaussian Grouping gg * 69.7 67.9 77.0 68.7 71.7 66.1
+OmniSeg3D omniseg3d 85.0 83.7 83.6 75.5 69.8 63.8
 Gaga (Ours) 92.3 90.8 72.0 63.3 71.2 68.4
-Table 1: Quantitative results for 3D query tasks on LERF-Mask.
-Gaga outperforms previous approaches, showcasing favorable performance in mIoU and BIoU with SAM.
-* denotes the results are reported in [32].
-Figure 5: Qualitative results on LERF-Mask.
-Our rendered segmentation exhibits fewer artifacts and delivers more accurate segmentation results than both prior 3D class-agnostic segmentation works and language embedding works.
-4.2 Open-vocabulary 3D Query on LERF-Mask
-We compare our method with 6 state-of-the-art methods on 3D scene understanding: LERF [14], LEGaussians [25], LangSplat [24], SA3D [3], Gaussian Grouping [32] and OmniSeg3D [33].
+* denotes results reported in gg .
+We compare our method with six state-of-the-art methods on 3D scene understanding: LERF lerf , LEGaussians legaussians , LangSplat langsplat , SA3D sa3d , Gaussian Grouping gg and OmniSeg3D omniseg3d .
 The first three methods focus on CLIP feature embedding.
 We calculate the relevancy between rendered CLIP features and query text features following their official implementation.
-The last three methods leverage 2D segmentation masks predicted by SAM [16].
-Since SAM does not support language prompts, SA3D, Gaussian Grouping and Gaga adopt Grounding DINO [19] to identify the mask ID in a 2D image and pick the corresponding 3D mask.
+The last three methods leverage 2D segmentation masks predicted by SAM sam .
+Since SAM does not support language prompts, SA3D, Gaussian Grouping and Gaga adopt Grounding DINO grounding_dino to identify the mask ID in a 2D image and pick the corresponding 3D mask.
 OmniSeg3D does not provide mask IDs, so we use Grounding DINO to identify a pixel that semantically aligns with the text prompt and select the corresponding 3D mask based on a similarity threshold of 0.9.
@@
 As shown in Fig.
-5, the visualization results of 3D query tasks with prompts “paper napkin” and “bag of cookies” further demonstrate the advancement of Gaga, as Gaga provides more clear segmentation masks without artifacts.
+5, the visualization results of 3D query tasks with prompts “paper napkin" and “bag of cookies" further demonstrate the advancement of Gaga, as Gaga provides more clear segmentation masks without artifacts.
+Figure 5: Qualitative results on LERF-Mask.
+Our rendered segmentation exhibits fewer artifacts and delivers more accurate segmentation results than both prior 3D class-agnostic segmentation works and language embedding works.
+Figure 6: Qualitative results on Replica and ScanNet.
+Gaga consistently predicts more accurate segmentation masks.
+Gaussian Grouping often covers the same object with different masks (row 1, 3, 5), creating large empty regions (row 1, 2, 3), or misidentifying similar instances (row 4, 5).
+4.3 3D Segmentation on Replica and ScanNet
+Table 2: Quantitative results on Replica and ScanNet.
+Gaga performs well with both 2D segmentation methods on two datasets, while the performance of Gaussian Grouping varies significantly with different segmentation methods.
 2D Seg. Replica ScanNet
 Model Method IoU Precision Recall IoU Precision Recall
-GG [32] EntitySeg 35.90 14.07 31.57 39.54 6.88 36.56
+GG (gg, ) EntitySeg 35.90 14.07 31.57 39.54 6.88 36.56
 Gaga (Ours) 41.45 59.74 45.59 43.08 35.46 49.73
-GG [32] SAM 21.76 25.00 19.72 34.24 18.70 32.61
+GG (gg, ) SAM 21.76 25.00 19.72 34.24 18.70 32.61
 Gaga (Ours) 46.21 39.53 50.62 45.06 22.88 51.02
-Table 2: Quantitative results on Replica and ScanNet.
-Gaga performs well with both 2D segmentation methods on two datasets.
-The performance of Gaussian Grouping (noted as GG) varies significantly with different 2D segmentation methods, whereas Gaga consistently delivers stable performance.
-Figure 6: Qualitative results on Replica and ScanNet.
-Gaga consistently predicts more accurate segmentation masks.
-Gaussian Grouping (noted as GG) often covers the same object with different masks (row 1, 3 and 5), creating large empty regions (row 1, 2 and 3), or misidentifying similar instances (row 4 and 5).
-2D Seg. Method EntitySeg SAM
-Model Training Data IoU $\uparrow$ IoU Drop $\downarrow$ IoU $\uparrow$ IoU Drop $\downarrow$
-GG [32] 30% 28.42 20.85 17.02 21.78
-Gaga (Ours) 40.40 6.22 42.53 5.61
-GG [32] 20% 24.56 31.35 16.02 26.38
-Gaga (Ours) 39.00 9.46 42.14 6.47
-GG [32] 10% 20.62 42.56 13.97 35.78
-Gaga (Ours) 35.92 16.60 39.06 13.30
-GG [32] 5% 10.00 72.15 6.77 68.87
-Gaga (Ours) 29.67 31.11 31.87 29.27
-Table 3: Quantitative results on Replica with limited data.
-Gaga consistently outperforms Gaussian Grouping with both 2D segmentation methods.
-The percentage of IoU drop shows that Gaga has greater robustness against reductions in training data.
-4.3 3D Segmentation on Replica and ScanNet
 Tab.
@@
 Row 1 and 2 depict the visualizations from the Replica dataset, while rows 3-5 showcase results from the ScanNet dataset.
-Gaussian Grouping [32] frequently assigns different mask IDs to the same object, resulting in inconsistent masks and empty regions.
+Gaussian Grouping gg frequently assigns different mask IDs to the same object, resulting in inconsistent masks and empty regions.
 Row 4 shows that Gaussian Grouping struggles to distinguish similar objects, whereas Gaga accurately identifies each object by leveraging 3D information.
+Figure 7: Qualitative results on Replica with limited data.
+The visualizations depict samples when using only 5% of the training data, where Gaga still produces high-quality segmentation.
+In contrast, Gaussian Grouping struggles to track objects accurately and leaves significant empty regions.
 4.4 3D Segmentation with Limited Data on Replica
-To demonstrate the robustness of Gaga against changes in training image quantity, we sparsely sample the Replica training set with ratios of 0.3, 0.2, 0.1, and 0.05.
-As depicted in Tab.
+Table 3: Results on Replica with limited data.
+Gaga outperforms Gaussian Grouping with both 2D segmentation methods.
+IoU drop shows that Gaga has greater robustness against reductions in training data.
+2D Seg. Method EntitySeg SAM
+Model Training Data IoU $\uparrow$ IoU Drop $\downarrow$ IoU $\uparrow$ IoU Drop $\downarrow$
+GG (gg, ) 30% 28.42 20.85 17.02 21.78
+Gaga (Ours) 40.40 6.22 42.53 5.61
+GG (gg, ) 20% 24.56 31.35 16.02 26.38
+Gaga (Ours) 39.00 9.46 42.14 6.47
+GG (gg, ) 10% 20.62 42.56 13.97 35.78
+Gaga (Ours) 35.92 16.60 39.06 13.30
+GG (gg, ) 5% 10.00 72.15 6.77 68.87
+Gaga (Ours) 29.67 31.11 31.87 29.27
+To demonstrate the robustness of Gaga against changes in training image quantity, we sparsely sample the Replica training set with different ratios.
+Depicted in Tab.
 3, Gaga consistently exhibits superior performance in terms of IoU, with approximately a 15% advantage using EntitySeg and a 25% advantage using SAM.
-Remarkably, when utilizing SAM, Gaga surpasses fully trained Gaussian Grouping with just 5% of the training data by more than 10% (31.87% vs.
-21.76%).
-We also compute the IoU drop compared to using all training images by:
-$IoU\ Drop(x\%)=\frac{IoU(100\%)-IoU(x\%)}{IoU(100\%)},$ (6)
-where $IoU(x\%)$ denotes the IoU achieved when $x\%$ of the training data is used.
-Compared to Gaussian Grouping, Gaga exhibits less sensitivity to decreases in the number of training images, as evidenced by smaller values in IoU drop.
+Remarkably, when utilizing SAM, Gaga surpasses fully trained Gaussian Grouping with just 5% of the training data by more than 10% (31.87% $v.s.$ 21.76%).
+We also compute the IoU drop compared to using all training images by: $IoU\ Drop(x\%)=\frac{IoU(100\%)-IoU(x\%)}{IoU(100\%)}.$ where $IoU(x\%)$ denotes the IoU achieved when $x\%$ of the training data is used.
+Gaga exhibits less sensitivity to decreases in the number of training images, as evidenced by smaller values in IoU drop.
 Visual results are shown in Fig.
@@
 With just 5% of the training data, Gaga delivers accurate segmentation masks, whereas Gaussian Grouping fails to provide masks for a significant portion of objects due to inaccurate tracking.
-Figure 7: Qualitative results on Replica with limited data.
-The visualizations depict samples when using only 5% of the training data.
-Even with limited data, Gaga consistently produces high-quality segmentation.
-In contrast, Gaussian Grouping struggles to track objects accurately and leaves significant empty regions.
 4.5 3D Segmentation on MipNeRF 360
@@
 We further showcase the performance of Gaga on the diverse MipNeRF 360 dataset.
-We provide visualization comparison on both indoor and outdoor scenes with Gaussian Grouping [32] in Fig.
-8, using SAM [16] (row 1, 2) and EntitySeg [23] (row 3, 4).
+We provide visualization comparison on both indoor and outdoor scenes with Gaussian Grouping gg in Fig.
+8, using SAM sam (row 1, 2) and EntitySeg entityseg (row 3, 4).
 We display two images for each scene to assess the consistency across different views.
-Gaga offers more detailed segmentation, while segmentation masks generated by Gaussian Grouping exhibit severe artifacts and empty regions.
-Additionally, inconsistency across two views exists in the rendering results of Gaussian Grouping, as shown in row 1 and row 3.
+Notably, inconsistency across the two views exists in the rendering results of Gaussian Grouping, as shown in row 1 and 3.
 4.6 Application: Scene Manipulation
-Figure 9: Scene manipulation results on MipNeRF 360 and Replica.
+Figure 9: Scene manipulation results.
 Gaga accurately identifies the cushion part of the footstool, whereas Gaussian Grouping colors it entirely.
@@
 Figure 10: Visual comparison of mask association methods.
-Gaga with 3D-aware memory bank achieves a superior visual quality and closer to the ground truth.
-Notice that the Video Tracker baseline mislabels the wall and floor, Our Memory Bank (w.
-All Gaussians) mislabels the floor.
-Gaga achieves high-quality and multi-view consistent 3D segmentation, beneficial for tasks like scene manipulation.
-We demonstrate the effectiveness of this application on the MipNeRF 360 [1] and Replica [27] datasets (see Fig.
-9).
-In the left image, we change the color of the footstool’s cushion and remove the stuffed animal on the armchair.
+Gaga with 3D-aware memory bank achieves a superior visual quality and is closer to the ground truth.
+Notice that the Video Tracker baseline mislabels the wall and floor.
+Our Memory Bank (w.
+All GS) mislabels the floor.
+Gaga achieves high-quality and multi-view consistent 3D segmentation, beneficial for scene manipulation tasks.
+We demonstrate the effectiveness of this application on the MipNeRF 360 mipnerf and Replica replica datasets in Fig.
+9.
+We change the color of the footstool’s cushion and remove the stuffed animal from the armchair.
 Gaga accurately identifies 3D Gaussians representing the cushion, while Gaussian Grouping fails, coloring the entire footstool maroon along with part of the sofa and some floating Gaussians.
@@
 4.7 Ablation Study on Mask Association Method
+Table 4: Ablation study. Our 3D-aware memory bank surpasses the previous video tracker baseline on all metrics.
 Baseline IoU Precision Recall
-SAM (Upper Bound) [16] 60.89 57.07 67.16
-Linear Assignment [26] 1.92 1.54 0.58
+SAM (Upper Bound) sam 60.89 57.07 67.16
+Linear Assignment panoptic_lifiting 1.92 1.54 0.58
 w/o. Mask Association 8.81 3.19 2.16
-Video Tracker [32] 21.76 25.00 19.72
-Our Memory Bank (All Gaussians) 42.30 40.48 46.51
+Video Tracker (gg, ) 21.76 25.00 19.72
+Our Memory Bank (All GS) 42.30 40.48 46.51
 Our Memory Bank 46.21 39.53 50.62
-Table 4: Ablation study on mask association methods. Our mask association method with the 3D-aware memory bank surpasses the previous video tracker baseline on all metrics.
 We conduct ablation studies to evaluate the effectiveness of the proposed mask association method on the Replica dataset.
-The baselines for comparison include: (1) Linear Assignment: Cost-based linear assignment mask association proposed in Panoptic Lifting [26].
+The baselines include: (1) Linear Assignment: Cost-based linear assignment mask association proposed in Panoptic Lifting panoptic_lifiting .
 (2) w/o.
 Mask Association: Lifting inconsistent 2D masks to 3D.
-(3) Video Tracker[32] is employed as a representative method.
+(3) Video Trackergg is employed as a representative method.
 (4) Our Memory Bank (w.
-All Gaussians): same as Gaga except that it selects all Gaussians splatted to the mask as its corresponding Gaussians.
+All GS): same as Gaga except that it selects all Gaussians splatted to the mask as its corresponding Gaussians.
 (5) Our Memory Bank: i.e., Gaga.
@@
 4 indicate that Gaga with the 3D-aware memory bank achieves superior performance compared to the previous method with video tracker.
-The cost-based linear assignment method used in [26] is not suitable for class-agnostic segmentation tasks, as it involves a significantly larger number of masks compared to panoptic segmentation.
+The cost-based linear assignment method used in panoptic_lifiting is not suitable for class-agnostic segmentation tasks, as it involves a significantly larger number of masks compared to panoptic segmentation.
 Comparison with Our Memory Bank (w.
 All Gaussians) baseline demonstrates the effectiveness of using depth maps to select corresponding 3D Gaussians for each mask.
-We also show the visual comparison in Fig.
+Visual comparison is shown in Fig.
 10.
-5 Conclusions
-We introduce Gaga, a framework that reconstructs and segments open-world 3D scenes using inconsistent 2D masks predicted by zero-shot segmentation models.
+5 Conclusions and Limitations
+Gaga reconstructs and segments open-world 3D scenes using inconsistent 2D masks predicted by zero-shot segmentation models.
 Gaga employs a 3D-aware memory bank to categorize 3D Gaussians and establishes mask association across different views by identifying the overlap between Gaussians projected to each mask.
 Results on various datasets show that Gaga outperforms previous methods with superior segmentation accuracy, multi-view consistency, and reduced artifacts.
-Additional application in scene manipulation further highlights Gaga’s high segmentation accuracy and practical utility.
+Although Gaga could leverage closed-set 2D panoptic segmentation masks to enable 3D panoptic segmentation, we leave this extension to future work beyond the scope of this paper.
 Supplementary Material
-1 Overview
+Appendix A Overview
 In this supplementary document, we provide further experimental results, including more results on scene manipulation and sparse view setting in Sec.
-3.
+B.
 We then delve into more experimental details of the datasets, metrics and implementation in Sec.
-4.
+C.
 More ablation studies are shown in Sec.
-5 and limitations are discussed in Sec.
-6.
-2 Supplementary Video
-Please watch the supplementary demo video for a comprehensive introduction and visual comparison between our method Gaga and the current state-of-the-art methods.
-The video features additional qualitative comparisons and an animation illustration of Gaga.
+D and limitations are discussed in Sec.
+E.
 Figure 11: Scene manipulation results on MipNeRF 360 and LERF-Mask.
 Gaga accurately identifies the flowerpot without affecting the color of the plant.
-Notice that Gaussian Grouping [32] creates a cyan region on the wooden door behind.
+Notice that Gaussian Grouping (gg, ) creates a cyan region on the wooden door behind.
 For the object removal and duplication tasks, Gaga can also provide more accurate results with fewer artifacts.
-3 Supplementary Experimental Results
-3.1 Results on Scene Manipulation
+Appendix B Supplementary Experimental Results
+B.1 Results on Scene Manipulation
 Gaga can accurately segment the Gaussians of a 3D object and edit their properties.
@@
 We provide additional results for the downstream scene manipulation task to further demonstrate the prospect of applying Gaga to real-world scenarios.
-On the “counter” scene of the MipNeRF 360 dataset [1], we change the color of the flowerpot to cyan and duplicate the glass jar.
-Gaussian Grouping [32] can not differentiate the plant and flowerpot, whereas Gaga generates a more accurate segmentation mask.
+On the “counter" scene of the MipNeRF 360 dataset mipnerf , we change the color of the flowerpot to cyan and duplicate the glass jar.
+Gaussian Grouping gg can not differentiate the plant and flowerpot, whereas Gaga generates a more accurate segmentation mask.
 Additionally, Gaga produces a clearer boundary and avoids artifacts on the iron tray when duplicating the glass jar.
-In the “figurines” scene of the LERF-Mask dataset [32], we transform the yellow duck to blue and remove the red toy chair.
+In the “figurines" scene of the LERF-Mask dataset gg , we transform the yellow duck to blue and remove the red toy chair.
 Gaga precisely changes only the duck’s color without affecting other objects, and achieves a more thorough removal of the red toy chair.
-3.2 Results on Sparsely Sampled Replica
+B.2 Results on Sparsely Sampled Replica
 Figure 12: Qualitative results on the sparsely sampled Replica.
@@
 Whereas Gaga exhibits a more robust performance against reductions in the number of images.
-4 Experimental Details
-4.1 Details on Datasets
-We employ the official script from Gaussian Splatting [13] for colmap to acquire camera poses and the initial point cloud.
+Appendix C Experimental Details
+C.1 Details on Datasets
+We employ the official script from Gaussian Splatting 3dgs for colmap to acquire camera poses and the initial point cloud.
 Consequently, the actual number of images utilized in the experiment might be lower than expected due to colmap process failures.
@@
 5 for the scene names used in the Replica and ScanNet datasets.
-LERF-Mask Dataset [32].
-LERF-Mask is based on the LERF dataset [14] and annotated with tasks and ground truth by the author of [32].
+LERF-Mask Dataset gg .
+LERF-Mask is based on the LERF dataset lerf and annotated with tasks and ground truth by the author of gg .
 It contains 3 scenes: figurines, ramen, and teatime.
-For each scene, 6-10 objects are selected as text queries, and Grounding DINO [19] is utilized to select the mask ID from the rendered segmentation.
-Replica Dataset [27].
-We select 8 scenes from the entire Replica Dataset the same as [35].
-We use the rendered results provided by authors of [35] and follow their data processing process: for each scene, we uniformly select 20% images as training data and 20% images as test data from all rendered RGB images.
+For each scene, 6-10 objects are selected as text queries, and Grounding DINO grounding_dino is utilized to select the mask ID from the rendered segmentation.
+Replica Dataset replica .
+We select 8 scenes from the entire Replica Dataset the same as semantic_nerf .
+We use the rendered results provided by authors of semantic_nerf and follow their data processing process: for each scene, we uniformly select 20% images as training data and 20% images as test data from all rendered RGB images.
 This results in 180 training images and 180 test images for each scene.
@@
 The number of test images remains at 180.
-ScanNet Dataset [10].
-DM-NeRF [29] selects 8 scenes from the entire ScanNet dataset.
+ScanNet Dataset scannet .
+DM-NeRF dm_nerf selects 8 scenes from the entire ScanNet dataset.
 Each scene has approximately 300 images for training and about 100 images for testing.
-We utilize 7 out of the 8 scenes, excluding ”scene 0024_00” due to the subpar 3D reconstruction results in both Gaussian Splatting [13] and Gaussian Grouping [32].
-MipNeRF 360 Dataset [1].
-We downsample the images by a factor of 4, consistent with the setting in [32], to accommodate the large size of the original images.
-For novel view synthesis evaluation, we set the sample step at 8, the same as the setting in [13].
+We utilize 7 out of the 8 scenes, excluding "scene 0024_00" due to the subpar 3D reconstruction results in both Gaussian Splatting 3dgs and Gaussian Grouping gg .
+MipNeRF 360 Dataset mipnerf .
+We downsample the images by a factor of 4, consistent with the setting in gg , to accommodate the large size of the original images.
+For novel view synthesis evaluation, we set the sample step at 8, the same as the setting in 3dgs .
+Table 5: Selected scenes in Replica and ScanNet. We select 8 scenes from the Replica dataset following (semantic_nerf, ), and 7 scenes from the ScanNet dataset following (dm_nerf, ).
 Dataset Scene Name
@@
 scene 0088_00 scene 0113_00 scene 0192_00
-Table 5: Selected scenes in Replica and ScanNet. We select 8 scenes from the Replica dataset following [35], and 7 scenes from the ScanNet dataset following [29].
-4.2 Details on Evaluation Metrics
+C.2 Details on Evaluation Metrics
 Given the disparate mask label assignments between the ground truth segmentation and the predicted segmentation for 3D objects, we find the best linear assignment between the labels based on IoU for quantitative evaluation.
@@
 $n_{gt}$leftarrow gt_masks.shape[1]
-   iou_matrix $\leftarrow$ torch.zeros(($n_{gt}$, max($n_{gt}$, $n_{pred}$)))
+[…]
 for $i$ in $n_{gt}$ do
@@
    for $k$ in $n_{image}$ do
-            iou_list.append(IoU(gt_masks[$k$][$i$], pred_masks[$k$][$j$]))
+[…]
    end for
-         iou_matrix[$i$][$j$] $\leftarrow$ iou_list.mean()
+   $i$][$j$] $\leftarrow$ iou_list.mean()
   end for
@@
 $gt\_indices$, $pred\_indices$leftarrow linear_assignment(iou_matrix)
-   paired_iou $\leftarrow$ iou_matrix[$gt\_indices$][$pred\_indices$]
+[…]
  iou $\leftarrow$ paired_iou.mean()
@@
 return iou, precision, recall
-4.3 Further Implementation Details
-For training vanilla 3D Gaussians, we maintain the same parameter setting as [13].
-To train the identity encoding, we freeze all the other attributes of Gaussians and use the same parameter setting as [32].
+C.3 Further Implementation Details
+For training vanilla 3D Gaussians, we maintain the same parameter setting as 3dgs .
+To train the identity encoding, we freeze all the other attributes of Gaussians and use the same parameter setting as gg .
 The identity encoding has 16 dimensions, and the rendered 2D identity encoding is in the shape of 16 $\times$ $h$$\times$ $w$, where $h$ and $w$ represent the height and width of the image.
@@
 All experiments are conducted on a single NVIDIA RTX 6000 Ada GPU.
+Table 6: Ablation study on the overlap threshold.
+If the overlap between the current mask and all groups in the memory bank falls below this threshold, we add this mask to the memory bank as a new group.
+Results indicate that the default setting of 0.1 generally yields better outcomes.
 figurines ramen teatime
@@
 0.3 85.7 84.0 45.6 43.3 71.6 69.0
-Table 6: Ablation study on the overlap threshold.
-If the overlap between the current mask and all groups in the memory bank falls below this threshold, we add this mask to the memory bank as a new group.
-Results indicate that the default setting of 0.1 generally yields better outcomes.
-5 Supplementary Ablation Study
-5.1 Ablation Study on Overlap Threshold
-We conduct additional ablation studies on the Gaussian overlap threshold using the Replica dataset [27], with SAM [16] as the 2D segmentation model.
+Appendix D Supplementary Ablation Study
+D.1 Ablation Study on Overlap Threshold
+We conduct additional ablation studies on the Gaussian overlap threshold using the Replica dataset replica , with SAM sam as the 2D segmentation model.
 During the group ID assignment process, if none of the existing groups in the memory bank exceed the overlap threshold with the current mask, we add the mask as a new group, indicating the discovery of a new 3D object.
@@
 To balance performance across all three metrics, we set the threshold to 0.1.
-5.2 Comparison on Mask Association Methods
+D.2 Comparison on Mask Association Methods
 Figure 13: Visual comparison between different mask association methods.
-Our 3D-aware memory bank offers more detailed associated masks, accurately tracks identical objects in the scene and assigns them different mask IDs.
-Conversely, video object tracker leaves empty regions in positions where it cannot track masks, and it struggles to provide consistent masks for the same object across views.
-We present visual comparison results for two mask association methods, video tracker [7] utilized by [32] and Gaga’s 3D-aware memory bank, in Fig.
+Gaga offers more detailed associated masks, accurately tracks identical objects in the scene and assigns them different mask IDs.
+Conversely, Gaussian Grouping leaves empty regions in positions where it cannot track masks, and it struggles to provide consistent masks for the same object across views.
+We present visual comparison results for two mask association methods, video tracker deva utilized by gg and Gaga’s 3D-aware memory bank, in Fig.
 13.
-In the “garden” scene of the MipNeRF 360 dataset, the video tracker struggles to track objects in the background, whereas Gaga provides associated results for each mask.
+In the “garden" scene of the MipNeRF 360 dataset, the video tracker struggles to track objects in the background, whereas Gaga provides associated results for each mask.
 For the scene in the ScanNet dataset, the video tracker fails to distinguish between four identical sofas, resulting in multiple masks for the same object.
@@
 In contrast, Gaga precisely locates each object, leading to improved mask association results and better pseudo labels for training segmentation features.
-6 Limitations
+Appendix E Limitations
 Though Gaga achieves SOTA performance compared to existing works, there are a few limitations and future works.